OpenStudy (anonymous):

Factor completely. -2k - k3 - 3k2 k(-k + 1)(k - 2) -k(k - 1)(k - 2) -k(k + 1)(k + 2)

OpenStudy (callisto):

What is the common factor of the three terms there?

OpenStudy (anonymous):

k

OpenStudy (callisto):

Hmm.. One more.. Can you see the negative signs?

OpenStudy (anonymous):

-k

OpenStudy (anonymous):

is it the 2nd one

OpenStudy (callisto):

Yup! So, take out the common factor and group the rest of the terms. You'll get this: \[-2k - k^3 - 3k^2\]\[=-k(2+ k^2 + 3k)\]Rearrange the terms a bit: \[ = -k (k^2+3k+2)\] From this, we can see that the 2nd one is nowhere near to the answer.